JAWAHARLAL NEHRU TECHNOLOGICAL UNIVERSITY HYDERABAD

JAWAHARLAL NEHRU TECHNOLOGICAL UNIVERSITY HYDERABAD B.Tech II PROBABILITY AND STATISTICS Previous Question Paper (December - 2018 - R15)

B.Tech | II | II-II | R15 | December - 2018

Answer PaperPaper Code: 124DKMaximum Marks: 75Time: 3 Hours

Note: This question paper contains two parts A and B. Part A is compulsory which carries 25 marks. Answer all questions in Part A. Part B consists of 5 Units. Answer any one full question from each unit. Each question carries 10 marks and may have a, b, c as sub questions.

Main

125 marks
a

The probability density function is

f(x) = 3(1-x^2) for 0 ≤ x ≤ 1, else where 0
Find K.

2 marks

Answer

For a valid probability density function, the integral over the entire range must equal 1. Given the function f(x) = k(1-x^2), we solve:

∫[0 to 1] k(1-x^2) dx = 1

Integrating gives k[x - x^3/3] from 0 to 1, resulting in k(1 - 1/3) = 1, which simplifies to 2k/3 = 1. Thus, k = 1.5.

b

Six cards are drawn from a pack of 52 cards. Getting a red card is a success. Find the probability of getting the success at least once.

3 marks

Answer

The probability of drawing a red card in a single draw is p = 26/52 = 0.5. For 6 cards, the number of red cards X follows a binomial distribution B(6, 0.5). The probability of at least one success is 1 - P(X=0).

P(X=0) = (6C0) * (0.5)^0 * (0.5)^6 = 1/64

Therefore, P(X ≥ 1) = 1 - 1/64 = 63/64 ≈ 0.9844.

c

The joint probability density function is given by

f(x,y) = (6-x-y)/8 for 0 < x < 2, 2 < y < 4, else where 0
Find the Marginal probability density function for X.

2 marks

Answer

Sign in to view the complete answer paper.

d

The equations of two Regression lines are y=11.64-0.5x, x=19.13-0.87y, find the Coefficient of Correlation.

3 marks

Answer

Sign in to view the complete answer paper.

e

Define tail I and tail II tests.

2 marks

Answer

Sign in to view the complete answer paper.

f

A Random sample of 1500 items is taken from a population whose standard deviation is 45. Find the maximum error.

3 marks

Answer

Sign in to view the complete answer paper.

g

Define Waiting time and Mean arrival rate.

2 marks

Answer

Sign in to view the complete answer paper.

h

What are the assumptions for the postulates for the poisson process?

3 marks

Answer

Sign in to view the complete answer paper.

i

Define an irreducible Marcov chain.

2 marks

Answer

Sign in to view the complete answer paper.

j

What are the properties of a transition matrix?

3 marks

Answer

Sign in to view the complete answer paper.

210 marks
a

A die is thrown 8 times. If getting 2 or 4 is a success, find the probability of getting: i) 4 successes. ii) P( x ≤ 3) (x being the number of failures)

5 marks

Answer

Sign in to view the complete answer paper.

b

Suppose that the weights of 800 male students are normally distributed with mean μ = 140 pounds with a standard deviation of 10 pounds. Find the number of students whose weight are: i) Between 138 and 148 ii) More than 152.

5 marks

Answer

Sign in to view the complete answer paper.

OR
310 marks
a

If the variance of a Poisson variate is 2, find the probability that: i) x =0 ii) 1< x < 4.

5 marks

Answer

Sign in to view the complete answer paper.

b

If the masses of 300 students are normally distributed with mean 68 kgs and standard deviation 3kgs. How many students have masses i) Greater than 72 kgs ii) Between 65 and 71 kgs.

5 marks

Answer

Sign in to view the complete answer paper.

4

The joint probability density function of the continuous random variables and is given by. Find f(x), f(y).

10 marks

Answer

Sign in to view the complete answer paper.

OR
5

Calculate the coefficient of rank correlation for the given data.

x68647550648075405564
y62586845816068485070
10 marks

Answer

Sign in to view the complete answer paper.

6

In a city A 20% of a random sample of 900 school boys had a certain slight physical defect. In another city B 18.5% of a random sample of 1600 school boys had the same defect. Is the difference between the proportions is significant at 0.05 level of significance.

10 marks

Answer

Sign in to view the complete answer paper.

OR
7

From the following data find whether there is any significant liking in the habbit of taking soft drinks among the categories of employees.

Soft drinksclerksteachersofficers
Pepsi102565
Thumps up153065
Sprite506030
10 marks

Answer

Sign in to view the complete answer paper.

8

The milk plant at a city distributes its products by trucks, loaded at the loading dock. It was its own fleet of trucks plus trucks of a private transport company. The trucks arrive at the interval of 20 minutes. The service time is 4 minutes.

10 marks
a

What is the probability that there are more than or equal to 4 trucks in the queue?

3 marks

Answer

Sign in to view the complete answer paper.

b

What is the waiting time of a truck in the queue?

3 marks

Answer

Sign in to view the complete answer paper.

c

What is the variance of queue length?

4 marks

Answer

Sign in to view the complete answer paper.

OR
9

Problems arrive at a computer centre in Poisson fashion with a mean arrival rate of 25 per hour. The average computing job requires 2 minutes of terminal time. Calculate the following:

10 marks
a

Average number of problems waiting for the queue.

3 marks

Answer

Sign in to view the complete answer paper.

b

The probability that queue length is greater than or equal to five.

3 marks

Answer

Sign in to view the complete answer paper.

c

The idle time.

4 marks

Answer

Sign in to view the complete answer paper.

10

There are three stores of groceries. There will be a shift from one to another. A study was made in January first and it was found that 1/4 was shopped at store1, 1/3 at 2 and 5/12 at 3. Every month store1 retain 90% of their customers and 10% transformed to store2. Store2 retain 5% of their customers and 85% transformed to store1 and store3. Store3 retain 40% of their customers and 50% transformed to store1 and 10% to store2.

10 marks
a

What proportion of customers retain by Feb 1st and March 1st

5 marks

Answer

Sign in to view the complete answer paper.

b

Find the long run proportions.

5 marks

Answer

Sign in to view the complete answer paper.

OR
11

In a presidential election at the end f June 40% of the voters registered for liberal, 45% for conservative and 15% for independent. Over one month, of the people those who registered for liberal 80% were retained, 15% changed to conservative and 5% to independent. Of the people those who registered for conservative 70% were retained, 20% changed to liberal and 10% to independent. Of the people those who registered for independent, 60% were retained, 20% changed to liberal and 20% to conservative.

10 marks
a

Write transition probability matrix

5 marks

Answer

Sign in to view the complete answer paper.

b

Find the % of the voters in each category at the end July

5 marks

Answer

Sign in to view the complete answer paper.

About this paper

This page contains the question paper of PROBABILITY AND STATISTICS from JAWAHARLAL NEHRU TECHNOLOGICAL UNIVERSITY HYDERABAD for B.Tech CEE II II-II, December - 2018, R15.

For the solved question paper with answers, switch to the Answer Paper tab.

Other papers in PROBABILITY AND STATISTICS

Loading Question Papers

Fetching data from database...